# Multistable Shape from Shading Emerges from Patch Diffusion

### v1

@@ 3.1 Dominant Global Lighting Constraint
Figure 3: Left: Illustration of multiscale sampling across two scales $x$ and $x^{\prime}$, with conditional images $c,c^{\prime}$ omitted for simplicity. The $N\&R$ subroutine injects noise to an earlier timestep $0<t<T$ and then resumes guided sampling (Fig. 2(b)) at that scale. Right: Optional intermediate guidance comes from lighting consistency, where each patch nominates a domininant light direction and then some patches flip in response to those nominations. Pseudocode is in the appendix.
Our guidance so far enforces global coherence, but even globally coherent surfaces can contain regions that independently undergo convex/concave flips without affecting their surround [29]. The top row of Fig. 5 provides a familiar example, where one can hide any three of the bumps/dents and then perceive the fourth as being either concave or convex. Yet, when one is allowed to examine the image as a whole and rotate it upside down, instead of perceiving $162^{4}=$ 16 interpretations, one sees only two: the bottom-right element is a bump (or dent) and the other three are its opposite. This behavior is explained by lighting. If one expects the dominant light direction to be similar everywhere on the surface, the four flips become tied together.
We can incorporate this notion of dominant light consistency into our model using an additional discrete guidance step, as depicted in the right of Fig. 3. This step can be applied to any global sample $\hat{x}_{0}$ and has three parts: (i) patches $\hat{x}_{0}^{u}$ independently nominate dominant light directions $\hat{l}^{u}$; (ii) we infer a single direction $\hat{l}$ that is most common among these nominations; and (iii) some patches perform a concave/convex flip to become more consistent with $\hat{l}$. Specifically, each patch $\hat{x}^{u}_{0}$ that is not too close to being planar (i.e., that has non-constant $\hat{x}^{u}_{0}(i)$) nominates its dominant light direction by computing the least-squares estimate according to shadowless Lambertian shading:
@@
We cluster the set $\{\hat{l}^{u}\}$ into two clusters using $k$-means, and we choose the center of the majority cluster as the dominant global direction $\hat{l}$. Then, each patch $u$ that belongs to the minority cluster performs a concave/convex flip $(p,q)\rightarrow(-p,-q)$ by multiplying the first two channels of $\hat{x}_{0}^{u}$ by $(-1)$. Finally, since the independent flips can cause discontinuities at patch seams, we always follow this discrete lighting guidance step by an inject Noise & Resume sampling (N&R) subroutine, where we add noise to an intermediate timestep via Eq. 1 and resume spatially-guided denoising from that timestep. Pseudocode is provided in Appendix A.1.
An advantage of this approach is that, unlike many previous computational approaches to shape from shading, it does not assume the lighting to be known a priori nor require that the lighting to be exactly spatially uniform across the surface. It also imposes no prior on the dominant light direction.
@@ 3.2 Multiscale Optimization
Since each local image patch can be explained by either concave or convex shapes, the terms in the spatial guidance energy (Eq. 7) are multimodal, and finding a global minimum is computationally difficult. Our experiments, like the one in Fig. 4, show that optimization at a single scale with random initial noise and gradient-descent guidance often gets trapped in poor local minima. To overcome this and also to utilize shading information from various spatial frequency and scales, we draw inspiration from work on Markov random fields [34] and introduce a multiscale optimization scheme. This is possible because our patch diffusion UNet and guidance can be applied to any resampled resolution $(sH)\times(sW)$ that is divisible by patch size $d$.
@@ 4 Experimental Results
Figure 4: Ablations, and comparison to human subjects using image and psychophysics data from [35]. Left: Ablations demonstrate the importance of each component. Right: Depth cross-sections extracted from four (integrated) samples from our full model exhibit relief-like variations similar to those reported across human subjects. (The dashed line is the depth used to render the input image.)
The input to our patch UNet has spatial dimension $d\times d$ with $d=16$ and $4$ channels, formed by the concatenation of $c^{u}$ and $x_{t}^{u}$. We train it using $d\times d$ patches from rendered images of the 3D objects in  [23], which are curated from Adobe Stock. We use Lambertian shading from random light directions, with a random albedo in [0.7,1.][0.7,1.][ 0.7 , 1 . ] and without cast shadows or global illumination effects. Our dataset contains around 8000 images ($256\times 256$) of 400 unique objects. Some examples are shown in the left of Fig. 2. The images contain occluding contours, and for empty background pixels $i$ we set $x_{0}(i)=(-1,-1,-1)$. We augment the training data by creating two convex/concave copies of each patch field $x^{u}_{0}$ that does not contain any background. At inference time, we use the DDIM sampler [45] with 50 sampling steps and with guidance. Additional details are in the appendix.
For comparison, we consider three deterministic approaches and two stochastic models. SIRFS [4] and Derender3D [50] are deterministic inverse-rendering models that estimate lighting and reflectance together with shape. They are among the few recent models that do not rely critically on having input occluding contour masks. Depth Anything [53] is a recent learning-based deterministic model for monocular depth estimation. It leverages a DINOv2 encoder [37] and a DPT decoder [41] and is trained for depth regression using 62M images. For comparisons to stochastic models, we include Marigold [24], which is derived from Stable Diffusion [42] and is fine-tuned for depth estimation. We also include Wonder3D [31], which also leverages a prior based on Stable Diffusion for improved generalization. Wonder3D is trained to generate consistent multi-view normal estimates on more than $30k$ 3D objects, and it achieves state-of-the-art results on 3D reconstruction benchmarks [10].
Figure 5: Normals produced by our model for various synthetic test surfaces rendered with directional light sources. For depth maps, brighter is closer. “Reference” depicts the mathematical ground truth shapes under directional light. Our reconstructions are more accurate and diverse than other methods.
@@ 4.1 Ablation Studies
Figure 4 analyses the key components of our model using a crop of a shape and image from the lab of James Todd [35] (the complete image is in the appendix). The left of the figure shows that when each patch is reconstructed independently, the resulting normals are inconsistent, because each patch may choose a different concave/convex mode as well as its various flattenings and tiltings. When spatial consistency guidance is applied at one scale, the global field is more consistent but suffers from discontinuous seams due to poor local minima. With multiscale sampling the seams improve, but separate bump/dent regions can still choose different modes without being consistent with any single dominant light direction. Finally, when lighting consistency is added, the output fields become more concentrated around two global modes—one that is globally convex (lit primarily from below) and another that is globally concave (lit primarily from above).
In the right of the figure, we compare samples from our model to depth profiles that were labeled by humans on the same image [35]. (These results have appeared previously in [29].) We generate four samples from our model’s globally-convex mode and integrate them to depth maps using [12]. Their cross-sections exhibit variations that are reminiscent of the variations between human subjects.
@@ 4.2 Ambiguous Images
Figure 6 visualizes distributions of shape reconstructions as 2D t-SNE plots (with perplexity equal to 30) by sampling 100 random seeds for our model and for Wonder3D. For reference, we also plot the t-SNE embeddings of a frontal plane, $\hat{x}_{0}(i)=(0,0,1)$, and of the two reference shapes. Our model covers both reference shapes whereas Wonder3D either covers only one or is close to a plane. These differences in coverage and accuracy are also apparent in terms of Wasserstein distance.
@@ 4.3 Real Images
Figure 7: Output samples using captured photographs. (a) For the plates image from [54], small regions (e.g., in the white box) can exhibit independent convex/concave flips between samples when lighting consistency is turned off; but when lighting consistency is turn on, only two global modes emerge. (b) Samples using photographs that we intentionally captured to appear ambiguous, especially when rotated by $180^{\circ}$. Half of the object in the first row was painted matte, leaving its other half glossy. Despite being trained entirely on synthetic data under idealized lighting, the model exhibits some generalization by producing plausible multistable outputs in real-world conditions.
Figure 8: Left: Output normals and integrated depth for an image taken from the web. Right: Results for photographs in the SfS dataset of [51], with median angular error from the ground truth normals (lower is better). Our model’s accuracy is on par with the best existing methods. Additional quantitative results are in the appendix.
We also evaluate our model using two categories of captured images. In each case, we resize the image to $256\times 256$ and use the multiscale schedule described in Appendix A.8.
Captured ambiguous images
Inspired by the plates image in [55], we captured a handful of images that induce multistable perceptions for human observers. We captured these images with a Sony $\alpha 7$SIII camera under lighting from area or point light sources, and some examples are in Fig. 7(b). We find that our model’s multimodality is qualitatively well aligned with perceptual multistability.
Shape from shading dataset [51]
This dataset contains diffuse objects captured with directional light sources and a dark background, along with ground truth surface normals for measuring accuracy. The right of Fig. 8 shows that our model produces normal maps that are on par with previous methods, even without knowledge of light source directions. The left of Figure 8 shows that our model can also produce a detailed and plausible shape estimates for an image taken from the web [2]. Additional quantitative results are in the appendix.
@@ 5 Related Work
Given the recent success of diffusion models in generating realistic images [19, 46, 25], many works have explored the power of patch-based diffusion, including for generating high-resolution panoramic images [3, 30]. Our method also leverages patch diffusion, but it departs from this work in two key ways. We do provide a global condition such as a text to each individual patch. Instead, each patch is conditioned only on the corresponding region of the input grayscale image. For this reason, we call it a bottom-up architecture. A second distinction is that, unlike [3, 30], we do not require an anchor patch or generate patches auto-regressively. Instead, we simultaneously guide all patches (e.g., $100$ patches for $160\times 160$ image) toward a coherent output using spatial consistency guidance.
Recently, Wang et al. [49] introduced a patch-based diffusion training framework that incorporates patch coordinates to reduce training time and storage cost. Patch-based diffusion has also been used for other tasks. Ozdenizci et al. [39] use overlapping patch diffusion to restore images in adverse weather, and Ding et al. [9] use it to synthesize higher resolution. All of these works use fairly large patches (e.g., $64\times 64$) and some of their components, such as feature-averaging or noise-averaging, are not appropriate for our shape from shading problem because of its inherent multi-modality. A convex sample and a concave sample cannot simply be averaged to improve the output. These challenges motivate the novel features of our model, including global consistency guidance and multiscale sampling.
@@ 6 Conclusion
A key limitation is our restriction to textureless and shadowless Lambertian shading. While this restriction is common in theoretical work  [27, 51, 17] and useful for creating ambiguous images, it is well-known that many ambiguities disappear in the presence of other cues such as glossy highlights, cast shadows, and repetitive texture. Also, since our model is predominantly bottom-up, it suffers when large regions of an image are covered by cast shadows (e.g., the shoulder region in Fig. 8). These types of regions often require non-local context like object recognition in order to be accurately completed. Incorporating more diverse materials and top-down signals into our model are important directions for future research.
@@ Acknowledgments
We thank Jianbo Shi for helpful discussions. We also thank James Todd, Benjamin Kunsberg, Steve Zucker and William Smith for kindly sharing their images and perceptual stimuli. This work was in part supported by JSPS 20H05951, 21H04893, and JST JPMJAP2305. It was also in part supported by the NSF cooperative agreement PHY-2019786 (an NSF AI Institute, http://iaifi.org).
@@ A.1 Algorithms pseudocode
We provide the pseudocode for the single-scale spatial consistency guided sampling (Alg. 1) and the lighting consistency guidance (Alg. 2). Here, we have hyperparameters $\lambda$ for weighting the smoothness and integrability loss, the $\eta_{t}$ as guidance update weight and $J_{t}$ as the number of noise update steps. In implementation, we set $\lambda=0.5$ and $J_{t}=3$. $\eta_{t}$ is resolution-dependent and is included with the scheduler specification in Appendix. A.8.
@@ A.2 Experimental setup details
We use a conditional UNet architecture similar to previous diffusion model [19] with input dimensions of 16 by 16 by 4 channels. The input to the UNet is a concatenation of the grayscale shading image, and a 3-channel normal map. We use the linear attention module [43] for better time and memory efficiency. The UNet consists of 4 downsampling and upsampling stages, with the commonly used ResNet [16] blocks, group normalization layers, attention layers and residual connections.
Dataset and Training Details
@@
We train the pixel-space conditional diffusion model on a dataset that we build from the UniPS dataset [23]. It contains about 8000 $256\times 256$ synthetic images of 400 unique objects from the Adobe3D Assets [1] rendered from different viewing directions. We render the objects with the shadow-less Lambertian model using the provided ground truth normal maps in [23], a randomly sampled directional light source within 60 degrees of the z-axis and an albedo value in [0.7,1.][0.7,1.][ 0.7 , 1 . ]. The normal map values outside of the object is masked to $(-1,-1,-1)$.
@@
We subdivide the image into non-overlapping $16\times 16$ patches and train our diffusion model to predict the noise at each sampled timestep using the smooth L1 loss. To train the diffusion UNet, we use the cosine variance schedule [36] with 300 timesteps. The model is trained with the AdamW optimizer and learning rate 2e-4 for 400 epochs, which takes about 40 hours on a single nVidia A100 GPU.
Data Augmentation
@@
During training, we augment the dataset with convex/concave flips of interior patches (those that do not include any occluding contour and background). In Sec.A.7, we compare two models trained with and without such augmentation. We show that while this augmentation leads to a more balanced multistable inference and thus smaller Wasserstein distance, a model trained naively from original image crops already gives multistable outputs on test images.
Sampling
@@
We use the DDIM sampler [45] with 50 sampling steps with guidance. Details of the multiscale sampling scheduler and guidance learning rate $\eta_{t}$ can be found in the appendix Sec.A.8.
Visualization and Evaluation Metric
@@
We visualize the output normal map from multiple samples to show the multistable reconstructions. We also show depth maps by integrating the normals using the method by Frankot & Chellappa [12]. To project the normal maps in lower dimensions, we first downsize them to 64 by 64 resolution and then use t-SNE plot (perplexity = 30) for visualization. We also compute the 1-Wasserstein distance from the 64 by 64 spatial resolution normal predictions.
Baseline models
@@
For testing using Wonder3D [31], we extract the frontal view prediction and use the default setting in their online demo as well as code base with a crop size of 256, and classifier-free guidance scale of 3 and with 50 sampling steps.

### v2

@@ 3.1 Dominant Global Lighting Constraint
Figure 3: Top: Illustration of multiscale sampling across two scales in a fine-coarse-fine “V-cycle”, with conditional images omitted for simplicity. In practice, our V-cycle covers more than two scales. Left: The $N\&R$ subroutine injects noise to an earlier timestep $0<t<T$ and then resumes guided sampling (Fig. 2(b)) at that scale. Right: Optional intermediate guidance comes from lighting consistency (LCG), where each patch nominates a dominant light direction and then some patches flip in response to those nominations. Pseudocode is in the appendix.
Our guidance so far enforces global coherence, but even globally coherent surfaces can contain regions that independently undergo convex/concave flips without affecting their surround [31]. The top row of Fig. 5 provides a familiar example. One can hide any three of the bumps/dents and then perceive the fourth as being either concave or convex. Yet, when one is allowed to examine the image as a whole and rotate it upside down, instead of perceiving $162^{4}=$ 16 interpretations, one sees only two: the bottom-right element is a bump (or dent) and the other three are its opposite. This behavior is explained by lighting. If one expects the dominant light direction to be similar everywhere on the surface, the four flips become tied together.
We can incorporate this notion of dominant light consistency into our model using an additional discrete guidance step, as depicted in the bottom right of Fig. 3. This step can be applied to any global sample $\hat{x}_{0}$ and has three parts: (i) patches $\hat{x}_{0}^{u}$ independently nominate dominant light directions $\hat{l}^{u}$; (ii) we identify a single direction $\hat{l}$ that is most common among these nominations; and (iii) some patches perform a concave/convex flip to become more consistent with $\hat{l}$.
Specifically, each patch $\hat{x}^{u}_{0}$ that is not too close to being planar (i.e., that has non-constant $\hat{x}^{u}_{0}(i)$) nominates its dominant light direction by computing the least-squares estimate according to shadowless Lambertian shading:
@@
We create two clusters in the set $\{\hat{l}^{u}\}$ using $k$-means and choose the center of the majority cluster as the dominant global direction $\hat{l}$. Each patch $u$ in the minority cluster undergoes a concave/convex flip $(p,q)\rightarrow(-p,-q)$ by multiplying $(-1)$ with the first two channels of $\hat{x}_{0}^{u}$. Since the independent flips can cause discontinuities at patch seams, we always follow this discrete lighting guidance step by an inject Noise & Resume sampling (N&R) subroutine, where we add noise to an intermediate timestep via Eq. 1 and resume spatially-guided denoising from that timestep. Pseudocode is provided in Appendix A.1.
This approach to lighting consistency has several advantages. Unlike many previous computational approaches to shape from shading, it does not assume the lighting to be known beforehand. Nor does it require the lighting to be exactly spatially uniform across the surface, which provides some resilience to global illumination effects. It imposes no prior on the dominant light direction (e.g., ‘lighting from above’), but one can imagine extending it to do so. And because our patch-based framework can be applied with or without lighting consistency guidance (see Appendix A.3), it may provide a mechanism in the future for modeling the way in which humans selectively enforce lighting consistency across an image [7, 40].
@@ 3.2 Multiscale Optimization
Since each local image patch can be explained by either concave or convex shapes, the terms in the spatial guidance energy (Eq. 8) are multimodal, and finding a global minimum is computationally difficult. Our experiments, like the one in Fig. 4, show that optimization at a single scale with random initial noise and gradient-descent guidance often gets trapped in poor local minima. To overcome this, and also to fully leverage shading information from various spatial frequencies and scales, we draw inspiration from work on Markov random fields [36] and introduce a multiscale optimization scheme. This is possible because our patch diffusion UNet and guidance can be applied to any resampled resolution $(sH)\times(sW)$ that is divisible by patch size $d$.
@@ 4 Experimental Results
Figure 4: Ablations, and comparison to human subjects using image and psychophysics data from [37]. Left: Ablations demonstrate the importance of each component. Right: Depth cross-sections extracted from four (integrated) samples from the convex mode of our full model exhibit relief-like variations similar to those reported across human subjects. (The dashed line is the depth that was used to render the input image.)
The input to our patch UNet is the concatenation of $c^{u}$ and $x_{t}^{u}$. It has $4$ channels and spatial dimension $d\times d$ with $d=16$. We train it using patches of size $d\times d$ extracted from rendered images of the 3D objects in [26] curated from Adobe Stock. We use Lambertian shading from random light directions, with a random albedo in $[0.5,1]$ and without cast shadows or global illumination effects. Our dataset contains around 8000 images ($256\times 256$) of 400 unique objects. Some examples are shown in the left of Fig. 2. The images contain occluding contours, and for empty background pixels $i,j$ we set $x_{0}(i,j)=(-1,-1,-1)$. We augment the training data by creating two convex/concave copies of each patch field $x^{u}_{0}$ that does not contain any background. At inference time, we use the DDIM sampler [48] with 50 sampling steps and with guidance. Additional details are in the appendix.
For comparison, we consider three deterministic approaches and two stochastic models. SIRFS [4] and Derender3D [53] are deterministic inverse-rendering models that estimate lighting and reflectance together with shape. They are among the few recent models that do not rely critically on having input occluding contour masks. Depth Anything [56] is a recent learning-based deterministic model for monocular depth estimation. It leverages a DINOv2 encoder [39] and a DPT decoder [44] and is trained for depth regression using 62M images. For comparisons to stochastic models, we include Marigold [27] which is derived from Stable Diffusion [45] and is fine-tuned for depth estimation. We also include Wonder3D [33], which likewise leverages a prior based on Stable Diffusion. Wonder3D is trained to generate consistent multi-view normal estimates on more than $30$k 3D objects, and it achieves state-of-the-art results on 3D reconstruction benchmarks [12].
Figure 5: Normals produced by our model for various synthetic test surfaces rendered with directional light sources. For depth maps, brighter is closer. “Reference” depicts the shapes—each with a convex/concave counterpart—that were used to render the input images. We find that our reconstructions are more accurate and diverse than other methods.
@@ 4.1 Ablation Studies
Figure 4 analyzes the key components of our model using a crop of a shape and image from the lab of James Todd [37] (the complete image is in Appendix A.5). The left of the figure shows that when each patch is reconstructed independently, the resulting normals are inconsistent, because each patch may choose a different concave/convex mode as well as its various flattenings and tiltings. When spatial consistency guidance is applied at one scale, the global field is more consistent but suffers from discontinuous seams due to poor local minima. With multiscale sampling the seams improve, but separate bump/dent regions can still choose different modes without being consistent with any single dominant light direction. Finally, when lighting consistency is added, the output fields become more concentrated around two global modes—one that is globally convex (lit primarily from below) and another that is globally concave (lit primarily from above).
In the right of the figure, we compare samples from our model to depth profiles that were labeled by humans on the same image [37]. (These results have appeared previously in [31].) We generate four samples from our model’s globally-convex mode and integrate them to depth maps using [15]. Their cross-sections exhibit variations with similar qualitative structure.
@@ 4.2 Ambiguous Images
Figure 6 visualizes distributions of shape reconstructions as 2D t-SNE plots (with perplexity equal to 30) by sampling 100 random seeds for our model and for Wonder3D. For reference, we also plot the t-SNE embeddings of a frontal plane, $\hat{x}_{0}(i,j)=(0,0,1)$ and of the two reference shapes. Our model covers both reference shapes whereas Wonder3D either covers only one or is close to a plane. These differences in coverage and accuracy are also apparent in terms of Wasserstein distance.
@@ 4.3 Real Images
Figure 7: Sampled reconstructions for real images. (a) For the ‘plates’ image from [57], regions such as the indicated box can exhibit independent convex/concave flips when lighting consistency is not used; but when lighting consistency is enforced, only two global modes emerge. (b) Sampled reconstructions for some multistable images we captured with illumination from a point or area light. (Rotate them by $180^{\circ}$ to enhance the alternative experience.) Note that half of the object in the first row was painted matte, and its other half was left glossy. Despite being trained entirely on synthetic data under idealized lighting, the model exhibits some generalization by producing plausible multistable outputs for these captured scenes.
Figure 8: Left: Reconstructed normals and integrated depth for an image taken from the web. Right: Reconstructions for images in the SfS dataset of [54] with median angular error from the ground truth normals (lower is better). Our model’s accuracy is on par with the best existing methods. Additional quantitative results are in the appendix.
We evaluate our model using a few different categories of captured images. In each case, we resize the image to $256\times 256$ (to accommodate the restrictions of some of the previous models), and we use the multiscale schedule described in Appendix A.10.
Captured ambiguous images.
Inspired by the ‘plates’ image in [58], we captured a handful of images that induce multistable perceptions for human observers. We captured these images with a Sony $\alpha 7$SIII camera under lighting from area or point light sources, and Fig. 7(b) shows some examples. We find that our model’s multimodality is qualitatively well aligned with perceptual multistability.
Shape from shading dataset [54].
This dataset contains diffuse objects captured with directional light sources and a dark background, along with ground truth surface normals for measuring accuracy. The right of Fig. 8 shows that our model produces normal maps that are on par with previous methods, even without knowledge of light source directions. Additional quantitative results are in Appendix A.8
Internet and astronomical images.
The left of Figure 8 shows that our model can produce a detailed and plausible shape estimate for a tone-mapped sRGB image taken from the web [2]. Appendix A.6 includes two satellite images of the surface of Mars and shows that our model reproduces the so-called crater illusion.
@@ 5 Related Work
Given the recent success of diffusion models in generating realistic images [22, 49, 28], many works have explored the power of patch-based diffusion, including for generating high-resolution panoramic images [3, 32]. Our method also leverages patch diffusion, but it departs from these work in two key ways. Unlike [3, 32], we do not generate patches auto-regressively or require an anchor patch. Instead, we simultaneously guide all patches (e.g., $100$ patches for a $160\times 160$ image) toward a coherent output using spatial consistency guidance. A second distinction is that we do not provide a global condition such as text to each individual patch. Instead, each patch is conditioned only on the corresponding crop of the input grayscale image, which is why we call it a bottom-up architecture. It shares the same spirit as previous work on inverse lighting [14], which also uses a bottom-up architecture to produce a variety of explanations that can then be integrated with top-down information.
Recently, Wang et al. [52] introduced a patch-based diffusion training framework that incorporates patch coordinates to reduce training time and storage cost. Patch-based diffusion has also been used for other tasks. Ozdenizci et al. [42] use overlapping patch diffusion to restore images in adverse weather, and Ding et al. [11] use it to synthesize images in higher resolution. All of these works use fairly large patches (e.g., $64\times 64$) and some of their components, such as feature-averaging or noise-averaging, are not appropriate for our shape from shading problem because of its inherent multi-modality. A convex sample and a concave sample cannot simply be averaged to improve the output. These challenges motivate the novel features of our model, including global consistency guidance and multiscale sampling.
@@ 6 Conclusion
A key limitation is our restriction to textureless and shadowless Lambertian shading. While this restriction is common in theoretical work [29, 54, 20] and useful for creating ambiguous images, it is well-known that many ambiguities disappear in the presence of other cues such as glossy highlights, cast shadows, and repetitive texture. Also, since our model is predominantly bottom-up, it suffers when large regions of an image are covered by cast shadows (e.g., the shoulder region in Fig. 8). These types of regions often require non-local context like object recognition in order to be accurately completed. Incorporating more diverse materials (e.g., as in [14]) and top-down signals into our model are important directions for future research.
@@ Acknowledgments
We thank Jianbo Shi for helpful discussions and Steven Zucker for suggesting the crater illusion. We also thank James Todd, Benjamin Kunsberg, Steven Zucker and William Smith for kindly sharing their images and perceptual stimuli. This work was in part supported by JSPS 20H05951, 21H04893, and JST JPMJAP2305. It was also in part supported by the NSF cooperative agreement PHY-2019786 (an NSF AI Institute, http://iaifi.org).
@@ A.1 Algorithms pseudocode
We provide the pseudocode for the single-scale spatial consistency guided sampling (Alg. 1) and the lighting consistency guidance (Alg. 2). Here, we have hyperparameters $\lambda$ for weighting the smoothness and integrability loss, $\eta_{t}$ as guidance update weight and $J_{t}$ as the number of noise update steps. The results in our paper use $\lambda=0.5$ and $J_{t}=3$. The parameter $\eta_{t}$ is resolution-dependent and is included with the schedule specification in Appendix A.10.
@@ A.2 Experimental setup details
We use a conditional UNet architecture similar to [22] with input spatial dimensions of $16\times 16$ and 4 channels. The input to the UNet is a concatenation of the grayscale shading image and a 3-channel normal map. We use a linear attention module [46] for better time and memory efficiency. The UNet consists of 4 downsampling and upsampling stages composed of the commonly used ResNet [19] blocks, group normalization layers, attention layers, and residual connections.
Dataset and Training Details.
@@
We train the pixel-space conditional diffusion model on a dataset that we build from the UniPS dataset [26]. It contains about 8000 $256\times 256$ synthetic images of 400 unique objects from the Adobe3D Assets [1] rendered from different viewing directions. We render the objects with the shadow-less Lambertian model using the provided ground truth normal maps in [26], a randomly sampled directional light source within 60 degrees of the $z$-axis, and an albedo value in $[0.5,1]$. The surface normal values outside of the objects are set to $(-1,-1,-1)$.
@@
We subdivide the image into non-overlapping patches of size $16\times 16$ and train our model to predict the noise at each sampled timestep using a smooth L1 loss. To train the diffusion UNet, we use the cosine variance schedule [38] with 300 timesteps. The model is trained using the AdamW optimizer for 500 epochs with learning rate 2e-4. It takes about 40 hours using one Nvidia A100 GPU.
Data Augmentation.
@@
During training, we augment the dataset with convex/concave flips of interior patches, meaning those that do not include any occluding contour and background. In Sec. A.7, we compare two models trained with and without such augmentation. While the augmentation leads to a more balanced multimodal output distribution and thus smaller Wasserstein distance, our model trained without augmentation is already capable of producing multistable outputs on test images.
Sampling.
@@
We use the DDIM sampler [48] with 50 sampling steps with guidance. Details of the multiscale sampling schedule and guidance learning rate $\eta_{t}$ can be found in Appendix A.10.
Visualization and Evaluation Metric.
@@
We visualize the output normal map from multiple samples to show the multistable reconstructions. We also show depth maps by integrating the normals using the method by Frankot & Chellappa [15].
To produce the t-SNE visualizations in Figs 6 and 13, we draw 100 samples from each model, and then downsize the sampled normal maps to $64\times 64$ resolution for computational efficiency. Then we project the samples to a 2-dimensional space using t-SNE with perplexity value of 30. To compute the 1-Wasserstein distance, we set the ground truth distribution to be a uniform distribution over the two reference normal maps: the one used to render the input image, and its convex/concave flip. When computing the 1-Wasserstein distance of model outputs to the ground truth distribution, all normal maps are first downsampled to $64\times 64$ resolution.
Baseline models.
@@
For testing with Wonder3D [33], we extract the frontal view prediction and use the default setting in their online demo and code base with a crop size of 256 by 256, classifier-free guidance scale of 3, and 50 sampling steps.